What is the formula of a compound if a sample of the compound contains 0.221 mol X, 0.088 mol Y, and 0.133 mol Z? Can somebody please explain the steps to this? Thanks!!
Answer #1

Compound contains 0.221 molx X, 0.088 mol Y, 0.133 mol Z .

So percent of X present is 0.221/(0.221+0.088+0.133) = 0.5.

of Y = 0.088/0.442~=0.2

so of Z= 0.3.

So the compound is X5Y2Z3
Answer #2

Sample contains X = 0.221 moles

Y = 0.088 moles

Z =0.133 moles

simplest whole ratio for X = 0.221/0.088 = 2.511

Y = 0.088/0.088 = 1

Z = 0.133/0.088= 1.511

for simplest ratio multiply all by 2

X = 5

Y = 2

Z = 3

Formula = X5Y2Z3
Answer #3

simple the formula is X0.221Y0.088Z0.133

now for getting all in natural numbers we need to make the least molar component 1 , so Y=1 , so multiply the stoichometric constant by 11.36

formula is X2.5YZ1.5 , still simplifying it , by multiplying each constant by 2 , we get the formula as X5Y2Z3